Circular No. 7788

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 A1 (LINEAR) AND C/2002 A2 (LINEAR) Two objects of asteroidal appearance separated by < 1^{o} and having

similar, somewhat unusual motion were reported by LINEAR on Jan. 8. Further observations were made following placement of the objects on the NEO Confirmation Page, and prediscovery observations in Nov.–Dec. 2001 were identified at the Minor Planet Center in LINEAR and NEAT data. Orbit computations for the two objects gave very similar results (basically differing only in T) and showed that the objects were only 0.7 AU from Jupiter at discovery. Observations by T. B. Spahr and P. Berlind with the 1.2-m reflector at the F. L. Whipple Observatory on 2002 Jan. 11.4 UT showed that the objects had faint, narrow tails around p.a. 250^{o}. Independent observations by R. H. McNaught with the 1.0-m reflector at Siding Spring on Jan. 11.6 showed that the first object, now designated C/2002 A1, was slightly diffuse with a very faint 25′′ tail in p.a. 260^{o}, and that the second object, C/2002 A2, was almost stellar with a very narrow 20′′ tail in p.a. 250^{o}. The respective discovery observations and orbital elements from MPEC 2002-A62 (which also contains all the relevant astrometry and ephemerides) are given below. The 2001 Dec. 13 and 17 observations of C/2002 A1 originally appeared on MPS 45271 under the designation 2001 XG_{115}. The minimum distances from Jupiter were < 0.4 AU around 2001 July 9 and 18 for C/2002 A1 and C/2002 A2, respectively.

Comet 2002 UT α_{2000} δ_{2000} m_{2}
C/2002 A1 Jan. 8.32291 7^{h}07^{m}18.^{s}51 +29^{o}45′55′′.5 18.7
C/2002 A2  8.32291 7 04 14.25 +29 39 58.4 18.4

Epoch = 2001 Nov. 27.0 TT

T = 2001 Dec. 2.2052 TT  ω = 19.^{o}1742
e = 0.743544  Ω = 82.0753 }2000.0
q = 4.698533 AU  i = 14.1563

a = 18.321040 AU n^{o} = 0.0125684 P = 78.420 years

Epoch = 2001 Nov. 27.0 TT

T = 2001 Dec. 9.0143 TT  ω = 19.^{o}3367
e = 0.739211  Ω = 82.2686 }2000.0
q = 4.708179 AU  i = 14.2299

a = 18.053597 AU n^{o} = 0.0128487 P = 76.709 years

2002 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green